# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ Abstract
With the increasing demand for seamless connectivity and intelligent communication, the integration of artificial intelligence (AI) and communication for sixth-generation (6G) network is emerging as a revolutionary architecture. This synergy deeply embeds AI into various layers of the communication network, optimizing resource allocation, enhancing network efficiency, and providing reliable performance in complex and dynamic environments. This paper presents a comprehensive overview of AI and communication for 6G networks, emphasizing their foundational principles, inherent challenges, and future research opportunities. We commence with a retrospective analysis of AI and the evolution of large-scale AI models, underscoring their pivotal roles in shaping contemporary communication technologies. The discourse then transitions to a detailed exposition of the envisioned integration of AI within 6G networks, delineated across three progressive developmental stages. The initial stage, AI for Network, focuses on employing AI to augment network performance, optimize efficiency, and enhance user service experiences. The subsequent stage, Network for AI, highlights the role of the network in facilitating and buttressing AI operations and presents key enabling technologies, including digital twins for AI and semantic communication. In the final stage, AI as a Service, it is anticipated that future 6G networks will innately provide AI functions as services and support application scenarios like immersive communication and intelligent industrial robots. Specifically, we have defined the quality of AI service, which refers to the measurement framework system of AI services within the network. In addition to these developmental stages, we thoroughly examine the standardization processes pertinent to AI in network contexts, highlighting key milestones and ongoing efforts. Finally, we outline promising future research opportunities that could drive the evolution and refinement of AI and communication for 6G, positioning them as a cornerstone of next-generation communication infrastructure.
@@ 1 Introduction
In recent years, the rapid development of wireless communication technology has profoundly reshaped various aspects of our society, driving unprecedented connectivity and enabling innovative applications. Following the widespread deployment and success of the fifth-generation (5G) networks, attention has shifted towards the sixth-generation (6G) wireless communication systems. With its enhanced capabilities, it is anticipated that 6G will bring transformative changes, including ultra-low latency, significantly higher data transmission rates, increased reliability, and ubiquitous connectivity. Among these advancements, integrating artificial intelligence (AI) into 6G networks is expected to be a game-changer, providing new paradigms and opportunities across multiple fields.
@@
AI technology has advanced rapidly over the past decade, particularly in machine learning, deep learning, and natural language processing. These advancements have enabled AI to play a crucial role across various industries. In the 5G era, AI has been successfully applied in aspects of wireless networks, such as network optimization, traffic prediction, and fault detection, significantly enhancing network performance and user experience. However, to truly achieve AI-native support in 6G networks, many unresolved issues still need to be addressed.
@@
The introduction of AI can enhance coding efficiency by utilizing compressed semantic information to transmit more data with less bandwidth, which helps alleviate network congestion and improve data transmission rates [1]. However, AI algorithms generate additional data that needs to be transmitted, such as model parameters, training data, and real-time feedback. This raises the question of whether the overall data volume in future networks will decrease or increase. This change in data volume will directly impact network design and architecture. Moreover, both base stations and user devices will employ AI algorithms for resource allocation to reduce energy consumption and improve resource utilization. However, the operation of AI itself may increase power consumption, leading to another question: Will energy consumption in future networks increase or decrease? In the pursuit of high efficiency, it is crucial to consider optimizing energy efficiency in AI applications to ensure the sustainable development of networks. Additionally, while networks can leverage AI algorithms to enhance transmission reliability and better respond to changing network conditions and user demands, the inherent uncertainty of AI algorithms—especially in complex and dynamic network environments—raises concerns about whether future networks will operate more reliably. These uncertainties may lead to decision-making errors, affecting user experience and overall network performance.
It is essential to consider integrating AI into the architecture, network elements, and functional processes of the new 6G system from a closer and deeper perspective to address these issues. Specifically, we need to re-examine the network’s design philosophy to ensure a genuine synergy is formed between AI and the 6G network.
@@
This comprehensive review explores the intricate relationship between 6G and AI, delving into the fundamental principles, key technologies, and potential applications this integration brings forth. We discuss the critical technical enablers of 6G, including advanced wireless communication techniques, spectrum management, and network architectures. Additionally, we examine the role of AI in optimizing network operations, enhancing security, and enabling intelligent decision-making processes. The review also highlights the open challenges to fully realizing the potential of 6G and AI integration. These challenges include data privacy, algorithmic transparency, energy efficiency, and the need for standardized frameworks to ensure interoperability and scalability. This review can be a valuable resource for researchers, practitioners, and policymakers involved in developing and deploying 6G and AI technologies by providing a holistic overview of the state-of-the-art and future directions. Through this collaborative effort, we can harness the full potential of these cutting-edge technologies to build a brighter, more connected, and more sustainable future.
@@
The organization of this paper is illustrated in Figure 1, and we summarize the main abbreviations used throughout this work in Table 1.
@@ 1.1 Requirements and Challenges in the Post-5G Era
1.1.1 More Efficient Resource Utilization
The rapid development of information technology and networks has led to a new round of business innovation and upgrading of service demand. The networks face the dual challenges of explosive growth in data traffic and the demand for massive devices to connect everything. According to Cisco statistics [2], global mobile data traffic grew more than tenfold from 2014 to 2019, and the number of connected devices worldwide had reached approximately 29.3 billion by the end of 2023. In the face of such an enormous traffic demand, it is of particular significance to explore how to utilize the limited network resources as fully as possible [3]. However, existing wireless communication systems still adopt a static spectrum management model, where authorized users have exclusive access. This model struggles to flexibly adapt to the significant spatiotemporal variations in traffic and the mobility characteristics of users, resulting in low spectrum resource utilization, as well as a huge power consumption of the base stations. Consequently, in the event that the network is capable of, by virtue of the intelligent prediction and dynamic adjustment capabilities furnished by AI, allocating network resources such as spectrum and power in a flexible manner in accordance with the real-time demands of various services, it will be feasible to curtail resource wastage or the scarcity of resources for specific services, thereby enhancing network performance and user experience.
1.1.2 More Intelligent Network Operation, Maintenance, and Management
Currently, the operation and maintenance management of mobile communication networks primarily rely on manual on-site troubleshooting, resulting in low levels of automation and high maintenance costs.
Table 1: List of Abbreviations
AI Artificial Intelligence MEC Multi-access Edge Computing
AI4NET AI for Network MARL Multi-agent Reinforcement Learning
AIaaS AI as a Service NLP Natural Language Processing
BS Base Station NTN Non-Terrestrial Networks
CNN Convolutional Neural Network NET4AI Network for AI
CP Control Plane NWDAF Network Data Analytics Function
DRL Deep Reinforcement Learning PFL Personalized Federated Learning
eMBB Enhanced Mobile Broadband PLMs Pre-trained Language Models
ENI Experiential Networked Intelligence QoAIS Quality of AI Service
FL Federated Learning QoS Quality of Service
GRU Gated Recurrent Unit RNN Recurrent Neural Networks
GAN Generative Adversarial Networks RF Random Forest
GPU Graphics Processing Units ResNet Residual Network
HARQ Hybrid Automatic Repeat Quest RSS Received Signal Strength
IoT Internet of Things SMPC Secure Multi-Party Computation
ITU International Telecommunication Union SLA Service Level Agreement
IMT International Mobile Telecommunications SA Service and Aspects
INC In-network Computing SNR Signal-to-Noise Ratio
KPI Key Performance Indicator UPF User Plane Function
LLMs Large Language Models ML Machine Learning
LSTM Long Short-Term Memory WG Working Group
uRLLC Ultra-Reliable Low Latency Communications mMTC Massive Machine Type Communications
@@ 1.1.2 More Intelligent Network Operation, Maintenance, and Management
This approach imposes high demands on the workload and expertise of maintenance personnel. The work mode is predominantly reactive, addressing issues as they arise rather than proactively preventing them. Consequently, providing rapid response reports and swift emergency handling capabilities is challenging. At the same time, 5G networks introduce new technologies, such as virtualization and mobile edge computing [4]. Edge sites are rapidly expanding, characterized by their large number, wide distribution, and heterogeneity. As the number of edge sites increases, maintenance teams must expand accordingly, leading to uncontrolled labor costs [5].
@@
Compared to traditional core network equipment’s typical millisecond-level fault detection efficiency, network resource virtualization introduces new potential fault points. Moreover, it is difficult for operators to delineate the responsibility boundaries among suppliers when faults occur. Using software mechanisms for fault detection, the virtual network function at the edge site typically experiences longer response times, which complicates timely network maintenance and management. Furthermore, the continuous emergence of new applications and services demands higher bandwidth, lower latency, more excellent reliability, and enhanced security from the network. Traditional network operation and maintenance management methods will find it increasingly difficult to adapt to the future’s more complex and dynamic network environments. The integration of AI and wireless networks can support the in-depth analysis and mining of a large amount of network data, achieve accurate fault prediction and rapid diagnosis, significantly reduce the relatively long network fault recovery time, and enhance network availability.
1.1.3 More Diverse and Heterogeneous Networks
The frequency bands for 6G networks will expand from the current 700 MHz to 2.6 GHz range, including millimeter waves, terahertz, and visible light frequencies. Coverage will extend from ground-based to three-dimensional, achieving integrated networking across air, space, and terrestrial domains. However, due to the dynamic changes in the network topology and different operating environments of non-terrestrial networks, the networking technologies used for terrestrial networks cannot be directly applied to non-terrestrial scenarios. It is necessary to research new networking technologies for integrated air-space-ground networks and consider new three-dimensional networking models [6]. On the other hand, with the extensive deployment of distributed edge computing and intelligent network nodes, resources such as computing and storage are being down to the edge nodes. For 6G to support the diverse business needs of various scenarios and meet the differentiated demands of users, it will require both centralized management of cellular networks and distributed collaborative cloud-edge integrated networking.
@@ 1.1.4 More Secure and Reliable Transmissions
From 2G to 5G, the design goal of mobile communication networks in terms of security and reliability has been to ensure the authenticity of user and network identities, prevent data interception and tampering during transmission, and primarily employ security mechanisms, such as network authentication, data encryption, and integrity checks, to safeguard communications. Moreover, these security measures have undergone several “add-on” enhancements and refinements. Establishing security measures is “post-hoc,” meaning that security mechanisms are added to the system after the design of the network communication functions is complete. Furthermore, some security vulnerabilities are only addressed in subsequent generations of mobile communication networks—for example, 5G remedied the capture attacks on user identity tags in 4G [7, 8]. The defense systems are deployed primarily based on a perimeter-centric defense architecture, most often erected on the premise of “known risks”. Protective devices, such as firewalls and intrusion detection systems, are positioned at the physical boundaries of mobile communication networks. Internally, these systems adopt a trusting stance, while externally, they use a ”patching” approach to ward off security risks.
Additionally, the currently released 5G network security protocol standards target primarily the Enhanced Mobile Broadband (eMBB) scenario. The standardization progress of security protocols for two other typical cases—Massive Machine Type Communications (mMTC) and Ultra-Reliable Low Latency Communications (uRLLC)—lags. In the post-5G era, mobile communication networks will increasingly focus on empowering vertical industry services. They will do so by opening up specific network capabilities to third parties through upper-layer interfaces or by providing customized security services tailored to the differentiated needs of various industries [9]. This approach is intended to effectively adapt to the new characteristics of mobile communication networks, which are more open, feature heterogeneous integration, include many terminals, and exhibit diverse connection types.
@@
1.2.1 Ground Vision
As new-generation information and communication technologies and applications such as big data, cloud computing, the Internet of Things (IoT), and AI develop, and as these intertwine deeply with information, communication, data and technology (ICDT), the technology landscape transitions into an era of “digital twins (DT) and ubiquitous intelligence” in the 6G era [10]. 6G networks will build a closely interconnected network space by providing communication services that integrate the physical and virtual worlds, enabling seamless interaction between the human society, physical world, and virtual world. This integration will create new value through the digital world, realizing the promising vision of “6G changing the world.”
In 2030 and beyond, mobile communication application scenarios will exhibit entirely new characteristics within the context of the DT world and pervasive intelligence. These scenarios will support ubiquitous wireless connectivity, big data, and new technologies such as AI, giving rise to three major application areas: intelligent living, intelligent production, and intelligent society. These areas will encompass integrated air-space-ground networks, communication and sensing interconnection, and intelligent interaction. 6G networks will not be confined to providing communication functions alone. Following the major trend of integrating sensing, communication, computing, AI, and security, 6G will expand from traditional single-function capabilities to new network capabilities such as sensing, computing, AI, and security.
@@ 1.2.2 Usage Scenarios of 6G
In June 2023, the International Telecommunication Union Radio Sector (ITU-R) approved the “Framework and Overall Objectives Recommendation for the Development of IMT for 2030 and Beyond” (referred to as “Recommendation”) at the 44th ITU-R WP 5D meeting. This document outlines the development goals, typical scenarios, and capability indicators for 6G, providing essential guidance for subsequent 6G technology and standards research. IMT-2030 (6G) has defined six significant scenarios. As shown in Figure 2, building on the “iron triangle” of IMT-2020 (5G), IMT-2030 (6G) extends outward to form a hexagon. On the outermost circle of the hexagon, four design principles applicable to all scenarios are listed: sustainability, ubiquitous intelligence, security/privacy/resilience, and connecting the unconnected.
@@
AI and Communication: The AI and communication integration scenario will support distributed computing and AI-driven applications. It will enable unprecedented use cases by leveraging data collection, local or distributed computation offloading, and distributed training and inference of AI models across intelligent nodes. Typical use cases in this scenario include assisted autonomous driving, autonomous collaboration between medical devices, offloading intensive computations across devices and networks, creating and predicting DT, and assisting collaborative robots. This scenario requires support for high regional traffic capacity, user experience data rates, low latency, and high reliability tailored to specific use cases. In addition to communication aspects, this use scenario is expected to include a range of new functionalities integrating AI and computing capabilities into IMT-2030. These include data collection, preparation, and processing from diverse sources, distributed AI model training, model sharing, and distributed inference across IMT systems, as well as coordination and linking of computing resources.

### v3

@@ Abstract
With the growing demand for seamless connectivity and intelligent communication, the integration of artificial intelligence (AI) and sixth-generation (6G) communication networks has emerged as a transformative paradigm. By embedding AI capabilities across various network layers, this integration enables optimized resource allocation, improved efficiency, and enhanced system robust performance, particularly in intricate and dynamic environments. This paper presents a comprehensive overview of AI and communication for 6G networks, with a focus on emphasizing their foundational principles, inherent challenges, and future research opportunities. We first review the integration of AI and communications in the context of 6G, exploring the driving factors behind incorporating AI into wireless communications, as well as the vision for the convergence of AI and 6G. The discourse then transitions to a detailed exposition of the envisioned integration of AI within 6G networks, delineated across three progressive developmental stages. The first stage, AI for Network, focuses on employing AI to augment network performance, optimize efficiency, and enhance user service experiences. The second stage, Network for AI, highlights the role of the network in facilitating and buttressing AI operations and presents key enabling technologies, such as digital twins for AI and semantic communication. The final stage, AI as a Service, it is anticipated that future 6G networks will innately provide AI functions as services, supporting application scenarios like immersive communication and intelligent industrial robots. Specifically, we have defined the quality of AI service, which refers to a framework for measuring AI services within the network. We further summarize the standardization process of AI for wireless networks, highlighting key milestones and ongoing efforts. In addition, we conduct an in-depth analysis of the critical challenges faced by the integration of AI and communications in 6G. Finally, we outline promising future research opportunities that are expected to drive the development and refinement of AI and 6G communications, paving the way for the next-generation communication infrastructure.
@@ 1 Introduction
In recent years, the rapid development of wireless communication technology has profoundly reshaped various aspects of our society, driving unprecedented connectivity and enabling innovative applications [1]. Following the widespread deployment and success of the fifth-generation (5G) networks, attention has shifted towards the sixth-generation (6G) wireless communication systems. With its enhanced capabilities, it is anticipated that 6G will bring transformative changes, including ultra-low latency, significantly higher data transmission rates, increased reliability, and ubiquitous connectivity [2]. Among these advancements, integrating artificial intelligence (AI) into 6G networks is expected to be a game-changer, providing new paradigms and opportunities across multiple fields [3].
@@
AI technology has advanced rapidly over the past decade, particularly in machine learning (ML) [4], deep learning (DL) [5], and natural language processing (NLP) [6]. These advancements have enabled AI to play a crucial role across various industries. AI, by virtue of its powerful data analysis and learning capabilities, can excavate massive data in wireless communication networks, realizing intelligent management and optimization of the network. In the 5G era, AI has been successfully applied in aspects of wireless networks, such as network optimization, traffic prediction, and fault detection, significantly enhancing network performance and user experiences [7]. However, to truly achieve AI-native support in 6G networks, many unresolved issues need to be addressed.
@@
The introduction of AI can enhance coding efficiency by utilizing compressed semantic information (SI) to transmit more data with less bandwidth, which helps alleviate network congestion and improve data transmission rate [8]. However, AI algorithms generate additional data that needs to be transmitted, such as model parameters, training data, and real-time feedback. This raises the question of whether the overall data volume in future networks will decrease or increase. This change in data volume will directly impact network design and architecture. Moreover, both base stations (BS) and user devices will employ AI algorithms for resource allocation to reduce energy consumption and improve resource utilization. However, the operation of AI itself may increase power consumption, leading to another question: Will energy consumption in future networks increase or decrease? In the pursuit of high efficiency, it is crucial to consider optimizing energy efficiency in AI applications to ensure the sustainable development of networks. Additionally, while networks can leverage AI algorithms to enhance transmission reliability and better respond to changing network conditions and user demands, the inherent uncertainty of AI algorithms—especially in complex and dynamic network environments—raises concerns about whether future networks will operate more reliably. These uncertainties may lead to decision-making errors, affecting user experience and overall network performance.
@@
It is essential to consider integrating AI into the architecture, network elements, and functional processes of the new 6G system from a closer and deeper perspective to address these issues. Specifically, we need to re-examine the network’s design philosophy to ensure a genuine synergy is formed between AI and the 6G network. This comprehensive review explores the intricate relationship between 6G and AI, delving into the fundamental principles, key technologies, and potential applications this integration brings forth. We discuss the critical technical enablers of 6G, including advanced wireless communication techniques, spectrum management, and network architectures. Additionally, we examine the role of AI in optimizing network operations, enhancing security, and enabling intelligent decision-making processes. The review also highlights the open challenges to fully realizing the potential of 6G and AI integration. This review can be a valuable resource for researchers, practitioners, and policymakers involved in developing and deploying 6G and AI technologies by providing a holistic overview of the state-of-the-art and future directions. Through this collaborative effort, we can harness the full potential of these cutting-edge technologies to build a brighter, more connected, and more sustainable future.
@@
The organization of this paper is illustrated in Figure 1, and we summarize the main abbreviations used throughout this work in Appendix A.
@@ 1.1 Requirements and Challenges in the Post-5G Era
1.1.1 Better Utilization of Spectrum Resources
The rapid development of information technology and networks has led to a new round of business innovation and upgrading of service demand and the volume of network traffic data has reached unprecedented levels [9]. The networks face the dual challenges of explosive growth in data traffic and the demand for massive devices to connect everything. As an essential mobile communication resource, spectrum remains the most critical element in improving network capacity [10]. Whether the limited spectrum resources can be better managed is the key to whether the network can guarantee high-quality services for users. However, existing wireless communication systems still adopt a static spectrum management model, where authorized users have exclusive access. This model lacks the ability of dynamic adjustment and is unable to flexibly allocate resources like spectrum and power in accordance with the real-time requirements of services. As a result, it is liable to lead to resource waste or insufficient resources for some services, which will impact network performance and user experiences.
1.1.2 Lower-Carbon Wireless Coverage
With the widespread adoption and continuous development of 5G networks, issues related to energy consumption and costs are becoming increasingly prominent. The energy consumption of the 5G system is about three time as that of the fourth-generation (4G) system due to its wider bandwidth, more channels, more complex equipment architecture [11]. Additionally, due to the use of higher frequency bands in 5G technology, there is a further reduction in the coverage area per BS. To achieve the same coverage targets, the number of 5G BS required would be three to four times that of 4G, significantly increasing the cost of network deployment for 5G BS. Moreover, the current operational management systems of actual networks lack flexibility and intelligence. The operational status of BS cannot be dynamically adjusted in precise accordance with real-time changes in regional user traffic. During periods and in areas with low traffic demand, the BSs may maintain a high workload, leading to energy waste. The mobile communication industry in the forthcoming era is inevitably trending towards a more environmentally friendly and sustainable pathway, with continuous endeavors in the innovation of green energy-efficient technologies.
1.1.3 More Efficient and Cost-effective Network O&M
@@
Currently, the operation and maintenance (O&M) management of mobile communication networks primarily rely on manual on-site troubleshooting, resulting in low levels of automation and high maintenance costs. This approach imposes high demands on the workload and expertise of maintenance personnel. The work mode is predominantly reactive, addressing issues as they arise rather than proactively preventing them. Consequently, providing rapid response reports and swift emergency handling capabilities is challenging. At the same time, 5G networks introduce new technologies, such as virtualization and mobile edge computing (MEC) [12]. Edge sites are rapidly expanding, characterized by their large number, wide distribution, and heterogeneity. As the number of edge sites increases, maintenance teams must expand accordingly, leading to uncontrolled labor costs [13].
@@
Compared to conventional core network (CN) equipment’s typical millisecond-level fault detection efficiency, network resource virtualization introduces new potential fault points. Moreover, it is difficult for operators to delineate the responsibility boundaries among suppliers when faults occur. Using software mechanisms for fault detection, the virtual network function at the edge site typically experiences longer response times, which complicates timely network maintenance and management. Moreover, the incessantly emerging novel applications and services have presented elevated requisites with respect to network bandwidth, latency, reliability, and security. The conventional methodologies for network O&M management will encounter difficulties in buttressing the more intricate and mutable network milieu in the perspective. Consequently, wireless networks are obliged to probe into innovative intelligent approaches for network operation, maintenance, and management, so as to guarantee the high efficacy and promptness of network deployment, operation, and maintenance.
1.1.4 More Personalized and Customized On-demand Service Capability
In the post-5G era, the mobile communication network, as a crucial infrastructure for digital social transformation, is no longer confined to the business domain of conventional mobile communication. Instead, it places greater emphasis on fulfilling the vastly diverse new demands for digital transformation across a multitude of industries. Specifically, the application scenarios of mobile communication networks are rich and diverse, covering multiple fields such as mobile communication, mobile internet, internet of things (IoT), smart cities, and satellite internet. Different application scenarios have significant differences in requirements for network user experience, data rate, service latency, reliability, etc. The network needs to possess the ability to provide personalized and customized services to meet the user needs of different scenarios [14]. Therefore, mobile communication networks need to develop in a more intelligent and efficient direction, continuously enhancing the personalized and customized on-demand service capabilities to meet new scenarios, new services, and new demands, improving user experience, and optimizing resource utilization.
@@ 1.1.5 More Secure and Reliable Transmissions
From 2G to 5G, the design goal of mobile communication networks in terms of security and reliability has been to ensure the authenticity of user and network identities, prevent data interception and tampering during transmission, and primarily employ security mechanisms, such as network authentication, data encryption, and integrity checks, to safeguard communications. Moreover, these security measures have undergone several “add-on” enhancements and refinements. Establishing security measures is “post-hoc”, meaning that security mechanisms are added to the system after the design of the network communication functions is complete. Furthermore, some security vulnerabilities are only addressed in subsequent generations of mobile communication networks—for example, 5G remedied the capture attacks on user identity tags in 4G [15, 16]. The defense systems are deployed primarily based on a perimeter-centric defense architecture, most often erected on the premise of “known risks”. Protective devices, such as firewalls and intrusion detection systems, are positioned at the physical boundaries of mobile communication networks. Internally, these systems adopt a trusting stance, while externally, they use a “patching” approach to ward off security risks.
Additionally, the currently released 5G network security protocol standards target primarily the enhanced mobile broadband (eMBB) scenario. The standardization progress of security protocols for two other typical cases: massive machine type communications (mMTC) and ultra-reliable low latency communications (uRLLC)—lags. In the post-5G era, mobile communication networks will increasingly focus on empowering vertical industry services. They will do so by opening up specific network capabilities to third parties through upper-layer interfaces or by providing customized security services tailored to the differentiated needs of various industries [17]. This approach is intended to effectively adapt to the new characteristics of mobile communication networks, which are more open, feature heterogeneous integration, include many terminals, and exhibit diverse connection types.
@@
1.2.1 Grand Vision
As new-generation information and communication technologies and applications such as big data, cloud computing, the IoT, and AI develop, and as these intertwine deeply with information, communication, data and technology, the technology landscape transitions into an era of “digital twin (DT) and ubiquitous intelligence” in the 6G era [18]. 6G networks will build a closely interconnected network space by providing communication services that integrate the physical and virtual worlds, enabling seamless interaction between the human society, physical world, and virtual world. This integration will create new value through the digital world, realizing the promising vision of “6G changing the world.”
In 2030 and beyond, mobile communication application scenarios will exhibit entirely new characteristics within the context of the DT world and pervasive intelligence. These scenarios will support ubiquitous wireless connectivity, big data, and new technologies such as AI, giving rise to three major application areas: intelligent living, intelligent production, and intelligent society. These areas will encompass integrated air-space-ground-sea networks, communication and sensing interconnection, and intelligent interaction. 6G networks will not be confined to providing communication functions alone. Following the major trend of integrating sensing, communication, computing, AI, and security, 6G will expand from conventional single-function capabilities to new network capabilities such as sensing, computing, AI, and security.
@@ 1.2.2 Usage Scenarios of 6G
In June 2023, the international telecommunication union (ITU) radio sector (ITU-R) approved the “Framework and Overall Objectives Recommendation for the Development of international mobile communications (IMT) for 2030 and Beyond” (referred to as “Recommendation”) at the 44th ITU-R WP 5D meeting. This document outlines the development goals, typical scenarios, and capability indicators for 6G, providing essential guidance for subsequent 6G technology and standards research. IMT-2030 (6G) has defined six significant scenarios. As shown in Figure 2, building on the “iron triangle” of IMT-2020 (5G), IMT-2030 (6G) extends outward to form a hexagon. On the outermost circle of the hexagon, four design principles applicable to all scenarios are listed: sustainability, ubiquitous intelligence, security/privacy/resilience, and connecting the unconnected.
@@
AI and Communication: The AI and communication integration scenario will support distributed computing and AI-driven applications. It will enable unprecedented use cases by leveraging data collection, local or distributed computation offloading, and distributed training and inference of AI models across intelligent nodes. Typical use cases in this scenario include assisted autonomous driving, autonomous collaboration between medical devices, offloading intensive computations across devices and networks, creating and predicting DT, and assisting collaborative robots. This scenario requires support for high regional traffic capacity, user experience data rates, low latency, and high reliability tailored to specific use cases. In addition to communication aspects, this usage scenario is expected to include a range of new functionalities integrating AI and computing capabilities into IMT-2030. These include data collection, preparation, and processing from diverse sources, distributed AI model training, model sharing, and distributed inference across IMT systems, as well as coordination and linking of computing resources.